OpenStudy (anonymous):

Solve the system, using substitution. Write the solution as an ordered pair. x = 2y - 2 and 3x - y = 14

OpenStudy (anonymous):

An ordered pair is basically just a point written as (x,y) Do you know how to solve this with substitution?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Ok, so substitution means to replace x or y for what you know x or y equals. For example, the first equation is x = -2y - 2, here we know what x equals, so we can put that in for x in the second equation which is 3x -y= 14 3(-2y -2) -2 = 14 Now we can solve for y. Once we get y we can put that in for "y" in one of the equations to solve for x This making sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok, see if you can solve for y from the substituted equation which is: 3(-2y -2) -y = 14

OpenStudy (anonymous):

im not good at solving things i freak under pressure.

OpenStudy (anonymous):

Ninja, shouldn't the correct equation be 3(2y-2)-y = 14? Where is the -2y coming from? it should be positive.

OpenStudy (anonymous):

i think lamaguineapig is right

OpenStudy (anonymous):

Yes it should..sorry I meant to type a y there but accidentally typed 2. >.>

OpenStudy (anonymous):

3(2y-2)-y=14 6y-6-y=14 5y-6=14 -6 -6 _____________ 5y=8 im not 100% sure im doing this right

OpenStudy (anonymous):

Yup, your almost right, but the last step you want to add 6 to both sides, so it should look like this 5y-6=14 +6 +6 ---------- 5y = 20 y = 4 Now you can put 4 in for y in one of those two equations.and solve for x! Try doing that. I'll double check your work for you. :)

OpenStudy (anonymous):

thats where i messed up!

OpenStudy (anonymous):

Yup. See if you can find x now. :)

OpenStudy (anonymous):

6

OpenStudy (anonymous):

Yup. I'll put it in the first equation to check it x = 2y - 2 x = 2(4) - 2 x = 8 -2 x = 6 So now we can put x and y in (x,y) and we have our answer!

OpenStudy (anonymous):

Here's a graph of the two lines. You can see that they intercept at (6,4) so our solution is correct!